OpenStudy (anonymous):

How much solute is dissolved in 31 ml of a sodium hydroxide solution with a pH of 13.3576?

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

@matt101

OpenStudy (matt101):

If you have the pH, you can find the pOH by subtracting the pH from 14. The pOH is similar to pH, just in terms of [OH-] rather than [H+]. Then answer the following questions and you'll have the answer to yours! What would the pOH of the solution be? What would the corresponding [OH-] be? How much OH-, in moles, are there? How much NaOH (the solute), in grams, was originally dissolved?

OpenStudy (anonymous):

Right so the pOH is 0.6424 but what do you mean corresponding OH-?

OpenStudy (anonymous):

@matt101 I'm confused

OpenStudy (matt101):

How do you go from pH to [H+]? It's the same process to go from pOH to [OH-]

OpenStudy (anonymous):

Wouldn't I do -log? @matt101

OpenStudy (matt101):

That's going from [OH-] to pOH. To go from pOH to [OH-], you need to find 10^(-pOH).

OpenStudy (anonymous):

Okay so 10^(-0.6424)=0.228

OpenStudy (anonymous):

How would you find moles of that? (Im sorry but pH confuses me a bit)

OpenStudy (matt101):

What you've just found is that [OH-]=0.228 M. You're given volume of solution in the question, so using this information you can find moles of OH- in the solution!

OpenStudy (anonymous):

Is there a formula for doing so?

OpenStudy (matt101):

Do you remember the formula for finding concentration?

OpenStudy (anonymous):

Do you mean M1V1=M2V2?

OpenStudy (matt101):

Sort of - that's for a dilution when you're dealing with a couple different concentrations. I'm talking about just one half of that equation: MV=n, or M=n/V, where M is the concentration, n is the moles of solute, and V is the volume of solution.

OpenStudy (anonymous):

Oh alright, so we are looking for volume of the solute? so can we just plug in the values we found?

OpenStudy (matt101):

Yup! And we're finding moles, because we have concentration and volume already.

OpenStudy (anonymous):

Oh alright, so MV=n (0.228M)(31mL)=7.068 And thats the answer?

OpenStudy (matt101):

Well remember the equation using volume in L, not mL, so it'll be (0.228M)(0.031L)=0.007068 moles. That's how many moles of OH- are in solution, meaning that's how many moles of NaOH were dissolved (because each molecule of NaOH contains 1 OH-). If you need the answer in grams, it's a simple conversion from moles to mass!

OpenStudy (anonymous):

Could you please show me the conversion? I just want to make sure I understand

OpenStudy (anonymous):

Like just use stoich?

OpenStudy (anonymous):

Oh I know so multiply the 0.007068 by the molar mass which is 40 so the answer is 0.28272g

OpenStudy (matt101):

Absolutely right!